ONLY ANSWER E,F,G2. Consider the convertible bond by Miser Electronics: par value1, 000 market price of convertible bond- $900 estimated strai
Answer #1

Solution=

a) Conversion value= Market price per share* no of shares

=$25*30 shares=$750

b)Market conversion price= Market priceof the bond / no of shares

=$900/30 shares

=$30

c)Conversion premium per share =(market price of the bond - conversion value)/no of shares

=(900-750)/30 shares

=$5 per share

e) Premiumover straight value={($750-$700)/$750}*100

=7.14%

f) income from bond=8.5%*1000=$85

Dividend from equity shares=$1 per share

Total no of share=25

Then dividdend =$25

Yield advantage of the bond=$85-$25=$60

g) premium pay back period = (bond Price- cinversion value)/ Yield advange of the bond

( 900-750)/60 =2.5 yera